OpenStudy (anonymous):

write next term of the sequence, then write rule. 2,6,12,20

hero (hero):

2 + 4 = 6 6 + 6 = 12 12 + 8 = 20 20 + ? =

OpenStudy (anonymous):

could be geometric try it out

OpenStudy (anonymous):

or do heros lol

OpenStudy (anonymous):

thank you guys! :)

hero (hero):

@Amberramirez, You figured out the correct numbers AND the rule as well? You should let us know what you got so we can make sure you got it right.

OpenStudy (anonymous):

@Hero I just got the next term. I'm still trying to figure out the rule. I'm not sure how to figure that out since the common ratio (right?) isn't constant. The next term is 30

OpenStudy (anonymous):

No

OpenStudy (anonymous):

@Hero

hero (hero):

Nevermind, it's not arithmetic sequence. Hang on

hero (hero):

Okay, so here's what we have: n = order of n terms in the sequence r = the result of the rule n = {1,2,3,4,5} r = {2,6, 12, 20, 30} How do we get from 1 to 2 Then from 2 to 6 Then from 3 to 12 and so on?

hero (hero):

Okay, well, basically, the rule is \[r_n = n(n + 1)\]

hero (hero):

\[r_1 = 1(1 + 1) = 1(2)= 2\]\[r_2 = 2(2 + 1) = 2(3)= 6\]\[r_3 = 3(3 + 1) = 3(4) = 12\] And so on

OpenStudy (anonymous):

oh okay that makes sense. thank you :)